OpenStudy (anonymous):

Your class sells 64 tickets to a play. A student ticket cost $1 and an adult ticket cost $2.50. If you collect $109 in total ticket sells, how many adult tickets did you sell?

OpenStudy (anonymous):

s + a = 64 ==> s = 64 - a s + 2.5a = 109 now sub 64 - a in for s in the 2nd equation 64 - a + 2.5a = 109 -- combine like terms 64 + 1.5a = 109 -- subtract 64 from both sides 1.5a = 109 - 64 1.5a = 45 a = 45/1.5 a = 30 Your class sold 30 adult tickets